# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 1 Introduction
The task is to find a parameterized function $y=F(x;\theta)$ that maps ensemble of sources $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times D}$ to targets $\boldsymbol{Y}^{*}=\{\boldsymbol{y}^{*}_{j}\}_{j=1}^{N}\in\mathbb{R}^{N\times d}$ with probability weights $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$. Here, $\boldsymbol{\mu}=(\mu_{1},\ldots,\mu_{M})^{\top}$ and $\boldsymbol{\nu}=(\nu_{1},\ldots,\nu_{N})^{\top}$ satisfy $\boldsymbol{\mu}^{\top}\mathbbm{1}_{M}=\sum_{i=1}^{M}\mu_{i}=1,\boldsymbol{\nu}^{\top}\mathbbm{1}_{N}=\sum_{j=1}^{N}\nu_{j}=1$ and $10<\mu_{i},\nu_{j}<$ 1. Note that $M$ may not be equal to $N$, and the same applies to $D$ and $d$. The absence of a one-to-one correspondence between the source and target data samples makes classical supervised regression methods inapplicable.
The OT solution finds an optimal $\theta$ by minimizing a loss function $\mathcal{L}(\theta)$ between the image of the source data $\boldsymbol{Y}_{\theta}=F(\boldsymbol{X};\theta)$ and the target data $\boldsymbol{Y}^{*}$, that is,
$\min_{\theta}\mathcal{L}(C^{\boldsymbol{Y}_{\theta}\rightarrow\boldsymbol{Y}^{*}},\boldsymbol{\mu},\boldsymbol{\nu}),$ (1)
where the cost matrix $C_{ij}^{\boldsymbol{Y}_{\theta}\rightarrow\boldsymbol{Y}^{*}}=c(\boldsymbol{y}_{i}(\theta),\boldsymbol{y}_{j}^{*})$ and $c$ is a function of cost between $\boldsymbol{y}_{i}$ and $\boldsymbol{y}_{j}^{*}$. Throughout this study, we assume that $c(\boldsymbol{y},\boldsymbol{y}^{*})$ is twice-differentiable, for instance, the squared Euclidean distance $c(\boldsymbol{y},\boldsymbol{y}^{*})=\|\boldsymbol{y}-\boldsymbol{y}^{*}\|_{2}^{2}$. Several popular OT distances are candidates for the loss function, including the Wasserstein distance, the entropy-regularized OT (EOT) distance, and the Sinkhorn distance [24]; see Section 2.1 for a brief review. Each of them leads to a bi-level optimization problem; for example, the EOT distance $\text{OT}_{\epsilon}(C^{\boldsymbol{Y}_{\theta}\rightarrow\boldsymbol{Y}^{*}},\boldsymbol{\mu},\boldsymbol{\nu})$ leads to
$\min_{\theta}\,\min_{\begin{subarray}{c}\Pi\in\mathbb{R}^{M\times N}_{\geq 0},\\ \Pi\mathbbm{1}_{N}=\boldsymbol{\mu},\Pi^{\top}\mathbbm{1}_{M}=\boldsymbol{\nu}\end{subarray}}\sum_{i=1}^{M}\sum_{j=1}^{N}C_{ij}^{\boldsymbol{Y}_{\theta}\rightarrow\boldsymbol{Y}^{*}}\Pi_{ij}+\epsilon\text{KL}(\Pi,\boldsymbol{\mu}\otimes\boldsymbol{\nu}).$
A crucial component of the optimization process is computing the derivatives of the loss function with respect to $\theta$, and hence the derivatives of the OT distance with respect to data $\boldsymbol{Y}_{\theta}$. First-order optimization methods require the gradients of the OT distances. Danskin’s theorem provides analytical gradients for the EOT distance [5, 8, 15, 14], but it does not apply to the Sinkhorn distance. Hence, the generic implicit differentiation method [20] has been widely applied to OT distances [10, 35, 36, 11, 6]. However, first-order optimization methods often converge slowly.
@@
$H(\Pi):=\left[\begin{array}[]{ccccccccc}\text{diag}(\Pi\mathbbm{1}_{N})&\Pi\\ (\Pi)^{\top}&\text{diag}(\Pi^{\top}\mathbbm{1}_{M})\end{array}\right]\in\mathbb{R}^{(M+N)\times(M+N)},$ (2)
where $\Pi$ is the coupling matrix. We show that when $\Pi$ has positive entries, $H(\Pi)$ has zero as a simple eigenvalue, and its effective condition number (i.e., the ratio of the largest and smallest positive eigenvalues) has upper and lower bounds depending on the spectral gap of $\Pi^{\top}\Pi$. In particular, we construct an example showing that $H(\Pi)$ can be severely ill-conditioned with the smallest positive eigenvalue at the order of $O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small, or $O(\frac{1}{N})$ when $N$ is large. Thus, when solving a linear system with $H$, proper regularization is crucial.
@@
  • Analytical derivatives and spectral analysis. We derive analytical gradients with respect to the data $\boldsymbol{Y}$ for EOT and Sinkhorn distances and Hessian for the EOT distance in Section 3.23.3. The spectral analysis in Section 4 helps us understand and resolve the numerical instability issue via a proper regularization in Section 3.4.
@@ 1.1 Outline
This work is organized as follows. Section 2 reviews the various (OT) distances and the Sinkhorn algorithm. Section 3.2 is devoted to the analytical and numerical computation of the gradients and Hessian, leading to an algorithm with proper regularization. Then, we analyze the spectrum of the matrix $H(\Pi)$ in Section 4. In Section 5, we examine the efficiency and accuracy of Hessian computation using a benchmark example and compare the results with other approaches. Then we apply the proposed algorithm to applications in Section 6, including the parameter estimation for shuffled regression of mixed Gaussian and 3D Point Clouds Registration of MobilNet10 dataset.
@@ 2 Optimal Transport Loss and Sinkhorn Algorithm
Ideally, we could find $\theta^{*}$ by minimizing the optimal transport distance between the parameterized source data $\boldsymbol{Y}_{\theta}$ and the target data $\boldsymbol{Y}^{*}$. We will first review some classical results in computational optimal transport [24] in this section. In the section, we ignore $\theta$ in the notation and $C$ is the abbreviation of $C^{\boldsymbol{Y}\rightarrow\boldsymbol{Y}^{*}}$, unless noted otherwise.
@@ 2.1.1 Wasserstein-2 Metric
One popular choice is to use the Wasserstein-2 metric as the optimal transport loss, equivalently, $\mathcal{L}(C,\boldsymbol{\mu},\boldsymbol{\nu})=W_{2}^{2}(C,\boldsymbol{\mu},\boldsymbol{\nu})$. To calculate the Wasserstein-2 metric, one has to solve a constraint optimization problem,
$W_{2}^{2}(C,{\boldsymbol{\mu}},\boldsymbol{\nu}):=\min_{\begin{subarray}{c}\Pi\in\mathbb{R}^{M\times N}_{\geq 0},\\ \Pi\mathbbm{1}_{N}=\boldsymbol{\mu},\Pi^{\top}\mathbbm{1}_{M}=\boldsymbol{\nu}\end{subarray}}\sum_{i=1}^{M}\sum_{j=1}^{N}C_{ij}\Pi_{ij},\quad C_{ij}=c(\boldsymbol{y}_{i},\boldsymbol{y}_{j}^{*}),$ (3)
where $c(\boldsymbol{y}_{i},\boldsymbol{y}_{j}^{*})=\|\boldsymbol{y}_{i}-\boldsymbol{y}_{j}^{*}\|_{2}^{2}$ is the cost of transport, and the coupling matrix $\Pi\in[0,1]^{M\times N}$ is the transport plan from the parameterized source data $\boldsymbol{Y}$ to the target data $\boldsymbol{Y}^{*}$. To solve the constraint optimization via linear programming, the computational complexity is $O((N+M)NM\log(N+M))$ [24], which is very expensive when $N,M$ are large. To overcome this issue, one often regularizes the objective function. Common regularization includes the EOT distance, the Sinkhorn distance, which we briefly review below.
@@ 2.1.2 Entropy-regularized OT (EOT) Distance
$\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu}):=\min_{\begin{subarray}{c}\Pi\in\mathbb{R}^{M\times N}_{\geq 0},\\ \Pi\mathbbm{1}_{N}=\boldsymbol{\mu},\Pi^{\top}\mathbbm{1}_{M}=\boldsymbol{\nu}\end{subarray}}\sum_{i=1}^{M}\sum_{j=1}^{N}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\Pi,\boldsymbol{\mu}\otimes\boldsymbol{\nu}),$ (4)
where the relative entropy between the coupling matrix $\Pi$ and the outer product $\boldsymbol{\mu}\bigotimes\boldsymbol{\nu}$ is $\text{KL}(\Pi,\boldsymbol{\mu}\otimes\boldsymbol{\nu}):=\sum_{i=1}^{M}\sum_{j=1}^{N}\Pi_{ij}\log\frac{\Pi_{ij}}{\mu_{i}\nu_{j}}$. This regularization drastically simplifies the study of the dual problem and further leads to the Sinkhorn algorithm for its unique numerical solution [24]. As $\epsilon$ goes to 0, EOT converges to the Wasserstein-2 distance at the rate of $\epsilon$ [20].
@@ 2.1.3 Sinkhorn Distance
Another candidate for the regularized OT loss is called the Sinkhorn distance, $\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$,
$\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu}):=\sum_{ij}C_{ij}\Pi^{*}_{ij},\text{ with }\Pi^{*}=\underset{\begin{subarray}{c}\Pi\in\mathbb{R}^{M\times N}_{\geq 0},\\ \Pi\mathbbm{1}_{N}=\boldsymbol{\mu},\Pi^{\top}\mathbbm{1}_{M}=\boldsymbol{\nu}\end{subarray}}{\operatorname{arg}\operatorname{min}}\;\sum_{ij}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\Pi,\boldsymbol{\mu}\otimes\boldsymbol{\nu}).$ (5)
Sinkhorn distance eliminates the contribution of the entropy regularization term from $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ to the total loss $\mathcal{L}$ after the transport plan $\Pi^{*}$ has been obtained. It actually gives even better approximation results and converges to the Wasserstein-2 distance exponentially fast. More precisely, we have $\left|\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})-W_{2}^{2}(C,\boldsymbol{\mu},\boldsymbol{\nu})\right|\leq c\exp(-1/\epsilon)$ [20].
@@ 2.2 Sinkhorn Algorithm
From now on, we firstly choose the EOT cost as the loss function, i.e., $\mathcal{L}(C,\boldsymbol{\mu},\boldsymbol{\nu})=\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ though we will discuss the others later. The computation of this quantity, i.e. the constraint optimization (4) is solved by the well-known Sinkhorn algorithm. For the notation consistency and ease on the readers, we will review necessary details of this algorithm. One can account for the constraints by introducing two slack variables, known as the dual potentials, $\boldsymbol{f}\in\mathbb{R}^{M},\boldsymbol{g}\in\mathbb{R}^{N}$, for each marginal constraint of (4). Because $\sum_{i}\sum_{j}\mu_{i}\nu_{j}=\sum_{i}\mu_{i}\sum_{j}\nu_{j}=1$, the corresponding augmented Lagrangian is
$\begin{split}\mathcal{L}(C,\Pi,\boldsymbol{f},\boldsymbol{g})&=\sum_{ij}C_{ij}\Pi_{ij}+\epsilon\sum_{ij}\Pi_{ij}\left(\log\frac{\Pi_{ij}}{\mu_{i}\nu_{j}}-1\right)\\ &\quad+\epsilon-\sum_{i}f_{i}(\sum_{j}\Pi_{ij}-\mu_{i})-\sum_{j}g_{j}(\sum_{i}\Pi_{ij}-\nu_{j})\,.\end{split}$ (6)
The first-order optimality condition $\frac{\partial\mathcal{L}(C,\Pi,\boldsymbol{f},\boldsymbol{g})}{\partial\Pi_{ij}}=0$ yields the expression of the optimal coupling $\Pi^{*}$ is
$\Pi_{ij}^{*}=\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f^{*}_{i}+g^{*}_{j}}{\epsilon}\right),$ (7)
@@
where $f^{*}_{i}$ and $g^{*}_{j}$ are the optimal dual potentials. It ensures that the optimal coupling matrix $\Pi^{*}$ is entry-wise positive. An intuitive scheme is to alternatively rescale rows and columns of the Gibbs kernel to satisfy the marginal constraint, which is called Sinkhorn algorithm[7, 30, 29, 21]. Numerically, however, this computation becomes unstable when $\epsilon$ is small. The stable Sinkhorn iteration is thus performed in the log-domain [7, 24],
$\boldsymbol{f}^{(l+1)}=\epsilon\log\boldsymbol{\mu}-\epsilon\log\left(K\exp(\boldsymbol{g}^{(l)}/\epsilon)\right),\;\boldsymbol{g}^{(l+1)}=\epsilon\log\boldsymbol{\nu}-\epsilon\log\left(K^{\top}\exp(\boldsymbol{f}^{(l+1)}/\epsilon)\right)$
@@
with the initial vector to be $\boldsymbol{g}^{(0)}=\mathbb{0}_{N}$ and the Gibbs kernel $K=\exp\left(-\frac{C}{\epsilon}\right)$. As $l$ goes to +$\infty$+ ∞, both converge to $\boldsymbol{f}^{*}$ and $\boldsymbol{g}^{*}$. In practice, the iteration stops when the 1-norm of marginal violation is within the threshold value.
Then we get the EOT distance $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ in terms of the dual problem,
@@
$\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})=\mathcal{L}(C,\Pi^{*},\boldsymbol{f}^{*},\boldsymbol{g}^{*})$ $=\left[\begin{array}[]{ccccccccc}\boldsymbol{\mu}^{\top}&\boldsymbol{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\boldsymbol{f}^{*}\\ \boldsymbol{g}^{*}\end{array}\right].$ (11)
@@
Overall the computational complexity of EOT to achieve $\tau$-approximate of the unregularized OT problem is $O(N^{2}\log(N)\tau^{-3})$ when $M=N$ [2, 24], which is significant improvement to the linear programming of Wasserstein-2 metric.
Similarly, $\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ gives
$\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})=\sum_{ij}C_{ij}\Pi^{*}_{ij}=\sum_{ij}\mu_{i}\nu_{j}C_{ij}e^{\frac{-C_{ij}+f_{i}^{*}+g_{j}^{*}}{\epsilon}}\,.$ (12)

### v2

@@ 1 Introduction
The task is to find a parameterized function $y=F(x;\theta)$ that maps ensemble of sources $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times D}$ to targets $\mathbf{Y}^{*}=\{\mathbf{y}^{*}_{j}\}_{j=1}^{N}\in\mathbb{R}^{N\times d}$ with probability weights $\bm{\mu}$ and $\bm{\nu}$. Here, $\bm{\mu}=(\mu_{1},\ldots,\mu_{M})^{\top}$ and $\bm{\nu}=(\nu_{1},\ldots,\nu_{N})^{\top}$ satisfy $\bm{\mu}^{\top}\mathbb{1}_{M}=\sum_{i=1}^{M}\mu_{i}=1,\bm{\nu}^{\top}\mathbb{1}_{N}=\sum_{j=1}^{N}\nu_{j}=1$ and $10<\mu_{i},\nu_{j}<$ 1. Note that $M$ may not be equal to $N$, and the same applies to $D$ and $d$. The absence of a one-to-one correspondence between the source and target data samples makes classical supervised regression methods inapplicable.
The OT solution finds an optimal $\theta$ by minimizing a loss function $\mathcal{L}(\theta)$ between the image of the source data $\mathbf{Y}_{\theta}=F(\mathbf{X};\theta)$ and the target data $\mathbf{Y}^{*}$, that is,
$\min_{\theta}\mathcal{L}(\mathbf{C},\bm{\mu},\bm{\nu}),$ (1)
where the cost matrix $C_{ij}=c(\mathbf{y}_{i}(\theta),\mathbf{y}_{j}^{*})$ and $c$ is a function of cost between $\mathbf{y}_{i}$ and $\mathbf{y}_{j}^{*}$. Throughout this study, we assume that $c(\mathbf{y},\mathbf{y}^{*})$ is twice-differentiable, for instance, the squared Euclidean distance $c(\mathbf{y},\mathbf{y}^{*})=\|\mathbf{y}-\mathbf{y}^{*}\|_{2}^{2}$. Several popular OT distances are candidates for the loss function, including the Wasserstein distance, the entropy-regularized OT (EOT) distance, and the Sinkhorn distance [26]; see Section 2.1 for a brief review. Each of them leads to a bi-level optimization problem; for example, the EOT distance $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ leads to
$\min_{\theta}\,\,\min_{\bm{\Pi}\in\mathbf{U}(\bm{\mu},\bm{\nu})}\sum_{i=1}^{M}\sum_{j=1}^{N}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\bm{\Pi},\bm{\mu}\otimes\bm{\nu}),$
where the convex polytope $\mathbf{U}(\bm{\mu},\bm{\nu})=\{\bm{\Pi}\in\mathbb{R}^{M\times N}_{\geq 0}:\bm{\Pi}\mathbb{1}_{N}=\bm{\mu},\bm{\Pi}^{\top}\mathbb{1}_{M}=\bm{\nu}\}$ is the set of bounded matrices with $M+N$ equality constraints.
A crucial component of the optimization process is computing the derivatives of the loss function with respect to $\theta$, and hence the derivatives of the OT distance with respect to data $\mathbf{Y}_{\theta}$. First-order optimization methods require the gradients of the OT distances. Danskin’s theorem provides analytical gradients for the EOT distance [5, 10, 17, 16], but it does not apply to the Sinkhorn distance. Hence, the generic implicit differentiation method [22] has been widely applied to OT distances [12, 37, 38, 13, 7]. However, first-order optimization methods often converge slowly.
@@
$\mathbf{H}:=\left[\begin{array}[]{ccccccccc}\text{diag}(\bm{\Pi}\mathbb{1}_{N})&\bm{\Pi}\\ (\bm{\Pi})^{\top}&\text{diag}(\bm{\Pi}^{\top}\mathbb{1}_{M})\end{array}\right]\in\mathbb{R}^{(M+N)\times(M+N)},$ (2)
where $\bm{\Pi}$ is the coupling matrix. When $\bm{\Pi}^{*}$ is the optimal coupling matrix, we denote this matrix (2) as $\mathbf{H}^{*}$. We show that when $\bm{\Pi}$ has positive entries, $\mathbf{H}$ has zero as a simple eigenvalue, and its effective condition number (i.e., the ratio of the largest and smallest positive eigenvalues) has upper and lower bounds depending on the spectral gap of $\bm{\Pi}^{\top}\bm{\Pi}$. In particular, we construct an example showing that $\mathbf{H}$ can be severely ill-conditioned with the smallest positive eigenvalue at the order of $O(e^{-\frac{1}{\epsilon}})$ when $\epsilon$ is small, or $O(\frac{1}{N})$ when $N$ is large. Thus, when solving a linear system with $\mathbf{H}$, proper regularization is crucial.
@@
  • Analytical derivatives and spectral analysis. We derive analytical gradients with respect to the data $\mathbf{Y}$ for EOT and Sinkhorn distances and Hessian for the EOT distance in Section 3.23.3. The spectral analysis in Section 4 helps us understand and resolve the numerical instability issue via a proper regularization in Section 3.4.
@@ 1.1 Outline
This work is organized as follows. Section 2 reviews the various (OT) distances and the Sinkhorn algorithm. Section 3 is devoted to the analytical and numerical computation of the gradients and Hessian, leading to an algorithm with proper regularization. Then, we analyze the spectrum of the matrix $\mathbf{H}$ in Section 4. In Section 5, we examine the efficiency and accuracy of Hessian computation using a benchmark example and compare the results with other approaches. Then we apply the proposed SGD-Newton approach to applications in Section 6, including the parameter estimation for shuffled regression of mixed Gaussian and 3D Point Clouds Registration of MobilNet10 dataset.
@@ 2 Optimal Transport Loss and Sinkhorn Algorithm
Ideally, we could find $\theta^{*}$ by minimizing the optimal transport distance between the parameterized source data $\mathbf{Y}_{\theta}$ and the target data $\mathbf{Y}^{*}$. We will first review some classical results in computational optimal transport [26] in this section. In the section, we ignore $\theta$ in the notation unless noted otherwise.
@@ 2.1.1 Wasserstein-2 Metric
One popular choice is to use the Wasserstein-2 metric as the optimal transport loss, equivalently, $\mathcal{L}(\mathbf{C},\bm{\mu},\bm{\nu})=W_{2}^{2}(\mathbf{C},\bm{\mu},\bm{\nu})$. To calculate the Wasserstein-2 metric, one has to solve a constrained optimization problem,
$W_{2}^{2}(\mathbf{C},{\bm{\mu}},\bm{\nu}):=\min_{\bm{\Pi}\in\mathbf{U}(\bm{\mu},\bm{\nu})}\sum_{i=1}^{M}\sum_{j=1}^{N}C_{ij}\Pi_{ij},\quad C_{ij}=c(\mathbf{y}_{i},\mathbf{y}_{j}^{*}),$ (3)
where $c(\mathbf{y}_{i},\mathbf{y}_{j}^{*})=\|\mathbf{y}_{i}-\mathbf{y}_{j}^{*}\|_{2}^{2}$ is the cost of transport, and the coupling matrix $\bm{\Pi}\in[0,1]^{M\times N}$ is the transport plan from the parameterized source data $\mathbf{Y}$ to the target data $\mathbf{Y}^{*}$. To solve the constrained optimization via linear programming, the computational complexity is $O((N+M)NM\log(N+M))$[26], which is very expensive when $N,M$ are large. To overcome this issue, one often regularizes the objective function. Common regularization includes the EOT distance, the Sinkhorn distance, which we briefly review below.
@@ 2.1.2 Entropy-regularized OT (EOT) Distance
$\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu}):=\min_{\bm{\Pi}\in\mathbf{U}(\bm{\mu},\bm{\nu})}\sum_{i=1}^{M}\sum_{j=1}^{N}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\bm{\Pi},\bm{\mu}\otimes\bm{\nu}),$ (4)
where the relative entropy between the coupling matrix $\bm{\Pi}$ and the outer product $\bm{\mu}\bigotimes\bm{\nu}$ is $\text{KL}(\bm{\Pi},\bm{\mu}\otimes\bm{\nu}):=\sum_{i=1}^{M}\sum_{j=1}^{N}\big{(}\Pi_{ij}\log\frac{\Pi_{ij}}{\mu_{i}\nu_{j}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\Pi_{ij}+\mu_{i}\nu_{j}}\big{)}$ [15]. This regularization drastically simplifies the study of the dual problem and further leads to the Sinkhorn algorithm for its unique numerical solution [26]. As $\epsilon$ goes to 0, EOT converges to the Wasserstein-2 distance at the rate of $\epsilon$ [22].
@@ 2.1.3 Sinkhorn Distance
Another candidate for the regularized OT loss is called the Sinkhorn distance, $\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$,
$\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu}):=\sum_{ij}C_{ij}\Pi^{*}_{ij},\text{ with }\bm{\Pi}^{*}=\underset{\bm{\Pi}\in\mathbf{U}(\bm{\mu},\bm{\nu})}{\operatorname{arg}\operatorname{min}}\;\sum_{ij}C_{ij}\Pi_{ij}+\epsilon\text{KL}(\bm{\Pi},\bm{\mu}\otimes\bm{\nu}).$ (5)
Sinkhorn distance eliminates the contribution of the entropy regularization term from $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ to the total loss $\mathcal{L}$ after the transport plan $\bm{\Pi}^{*}$ has been obtained. It gives even better approximation results and converges to the Wasserstein-2 distance exponentially fast, that is, we have $\left|\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})-W_{2}^{2}(\mathbf{C},\bm{\mu},\bm{\nu})\right|\leq c\exp(-1/\epsilon)$ [22].
@@ 2.2 Sinkhorn Algorithm
From now on, we firstly choose the EOT cost as the loss function, i.e., $\mathcal{L}(\mathbf{C},\bm{\mu},\bm{\nu})=\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ though we will discuss the others later. The computation of this quantity, i.e., the constrained optimization (4) is solved by the well-known Sinkhorn algorithm. As a preliminary, we briefly review this algorithm. One can introduce two slack variables, known as the dual potentials, $\mathbf{f}\in\mathbb{R}^{M},\mathbf{g}\in\mathbb{R}^{N}$, for each marginal constraint of (4). Also because of the relative entropy term $\log\frac{\Pi_{ij}}{\mu_{i}\nu_{j}}$, $\Pi_{ij}$ will never be negative and we can drop non-negativity constraints $\Pi_{ij}\geq 0$. As a result, the corresponding Lagrangian is
$-ϵ\begin{split}L(\mathbf{C},\bm{\Pi},\mathbf{f},\mathbf{g})&=\sum_{ij}\left(C_{ij}\Pi_{ij}+\epsilon\Pi_{ij}\big{(}\log\frac{\Pi_{ij}}{\mu_{i}\nu_{j}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1}\big{)}-f_{i}(\Pi_{ij}-\mu_{i})-g_{j}(\Pi_{ij}-\nu_{j})\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\epsilon}\,.\end{split}$
The strong duality [6, 8] holds, $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})=\max_{\mathbf{f},\mathbf{g}}\min_{\bm{\Pi}}L(\mathbf{C},\bm{\Pi},\mathbf{f},\mathbf{g})$. Also, the first-order optimality conditions of the inner problem yield the expression of the optimal coupling matrix $\bm{\Pi}^{*}$ as the function of $\mathbf{f}$ and $\mathbf{g}$, $\Pi_{ij}^{*}=\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f_{i}+g_{j}}{\epsilon}\right)$. It ensures that the optimal coupling matrix $\bm{\Pi}^{*}$ is entry-wise positive. The first-order optimality conditions of the outer problem yield the marginal constraints,
$\sum_{j}\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f^{*}_{i}+g^{*}_{j}}{\epsilon}\right)=\mu_{i},\ \ \sum_{i}\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f^{*}_{i}+g^{*}_{j}}{\epsilon}\right)=\nu_{j}.$ (6)
@@
An intuitive scheme to solve these nonlinear equations (6) for optimal dual variable $\mathbf{f}^{*}$ and $\mathbf{g}^{*}$ is to alternatively rescale rows and columns of the Gibbs kernel to satisfy the marginal constraint, which is called Sinkhorn algorithm[9, 32, 31, 23]. Numerically, however, this computation becomes unstable when $\epsilon$ is small. The stable Sinkhorn iteration is thus performed in the log-domain [9, 26],
$\mathbf{f}^{(l+1)}=\epsilon\log\bm{\mu}-\epsilon\log\left(\mathbf{K}\exp(\mathbf{g}^{(l)}/\epsilon)\right),\;\mathbf{g}^{(l+1)}=\epsilon\log\bm{\nu}-\epsilon\log\left(\mathbf{K}^{\top}\exp(\mathbf{f}^{(l+1)}/\epsilon)\right)$
@@
with the initial vector to be $\mathbf{g}^{(0)}=\mathbb{0}_{N}$ and the Gibbs kernel $\mathbf{K}=\exp\left(-\frac{\mathbf{C}}{\epsilon}\right)$. As $l$ goes to +$\infty$+ ∞, both converge to $\mathbf{f}^{*}$ and $\mathbf{g}^{*}$. In practice, the iteration stops when the 1-norm of marginal violation is within the threshold value.
Finally we get the optimal coupling matrix $\bm{\Pi}^{*}$ and EOT distance $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$,
@@
$\Pi_{ij}^{*}=\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f^{*}_{i}+g^{*}_{j}}{\epsilon}\right),\ \ \text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})=\left[\begin{array}[]{ccccccccc}\bm{\mu}^{\top}&\bm{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbf{f}^{*}\\ \mathbf{g}^{*}\end{array}\right].$ (10)
@@
Overall the computational complexity of EOT to achieve $\tau$-approximate of the unregularized OT problem is $O(N^{2}\log(N)\tau^{-3})$ when $M=N$ [2, 26], which is a significant improvement to the linear programming of Wasserstein-2 metric.
Similarly, $\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ gives
$\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})=\sum_{ij}C_{ij}\Pi^{*}_{ij}=\sum_{ij}\mu_{i}\nu_{j}C_{ij}e^{\frac{-C_{ij}+f_{i}^{*}+g_{j}^{*}}{\epsilon}}\,.$ (11)